# Did Translation Models Get More Robust Without Anyone Even Noticing?

### v1

@@ 1 Introduction
For years, the conventional wisdom has been that neural machine translation (MT) models are highly sensitive to source-side artificial and natural noise at inference time (Belinkov and Bisk, 2018). This insight has motivated many works that seek to make MT models more robust to noise through either specialized training (Ebrahimi et al., 2018; Karpukhin et al., 2019; Park et al., 2020; Vaibhav et al., 2019) or bespoke architectures (Rust et al., 2022; Salesky et al., 2021). However, MT is increasingly being performed in a different paradigm than when these analyses and architectures were created. Previously, models were mostly trained from scratch on task-specific data, often for a single language pair. Nowadays, strong results often depend on large pretrained encoder-decoder models (NLLB Team et al., 2022), instruction-tuned large language models (LLMs) like TowerLLM (Alves et al., 2024), or opaque proprietary systems like ChatGPT.111https://chat.openai.com/ These huge models may make existing robustness techniques more expensive (specialized training takes longer with larger models) or impossible (specialized architectures cannot be grafted onto an existing pretrained system). So the question is, how necessary are these robustness techniques for LLMs? Are they still needed to mitigate the brittleness of subword-level transformers, or have larger models and training sets made today’s models sufficiently robust on their own?
In this work, we investigate these questions through experiments on social media text and synthetically noised corpora. These experiments have complementary roles: social media text contains diverse noise phenomena, but isolating their effect is not straightforward because the errors are not labeled. On the other hand, synthetic errors may differ in major ways from “naturally occurring” noise, but they are interpretable and controllable, offering a way to measure noise in vitro. By evaluating on a broad spectrum of error types, we can paint a more vivid picture of what kinds of noise, and at what quantities, cause problems for MT systems.
We make the following contributions:222Our code is available at https://github.com/deep-spin/robust-mt.
We show (§3) that large pretrained models are much more robust to synthetic source-side errors than conventional single-pair NMT models (see Figure 1), even when their performance is similar on clean data. These results hold across several language pairs and varieties of noise, even though the large models lack architectural features that obviously encourage robustness to character noise.
@@
We show (§5) that source correction pipelines can be an effective approach to mitigate the impact of synthetic noise without substantially worsening performance on clean data, although they are significantly less effective with stronger models, suggesting that the benefits of source correction and model robustness are not complementary. Source correction is less effective on social media data, likely because there are not enough errors to outweigh the risk of error propagation.
@@ 2.1 Architectures for MT
In recent years, mainstream MT techniques have been based on the transformer (Vaswani et al., 2017), which uses multi-headed self-attention to mix information across time steps. In the original work, transformers used an encoder-decoder paradigm similar to recurrent MT models (Bahdanau et al., 2014). These models pair an encoder over the source with a decoder, an autoregressive language model that predicts target tokens one at a time. These tokens usually come from a learned subword vocabulary (Kudo, 2018; Sennrich et al., 2016). Initially, transformer MT models were trained from scratch for a single language pair on parallel data from sources such as the OPUS parallel corpus collection (Tiedemann, 2012).
@@
Although single language pair models often perform well, they struggle in the absence of large quantities of data, making it difficult to achieve good results on low resource language pairs. This problem can be mitigated somewhat through multilingual training with systems like M2M-100 (Fan et al., 2021) and NLLB-200 (NLLB Team et al., 2022). Low resource language pairs often benefit from training data in other languages. One challenge is language imbalance – the subword vocabulary and training procedure need to be designed to allow strong performance across covered language pairs in spite of this imbalance.
@@
In parallel to these MT-centric developments, transformers have increasingly been used in a transfer learning set-up in which a large model is pretrained on some generic objective for which massive data is available. Such a model can then be finetuned on a particular downstream task or set of tasks. When the pretraining objective is language modeling (Radford et al., 2018), this makes it straightforward to use the model for text generation tasks such as MT. Recent generation-oriented models are often decoder-only: to use such a model for MT, the source is simply treated as part of the decoder’s context. Such models have shown some success for MT (Hendy et al., 2023). In recent times, the paradigm has shifted from traditional finetuning to instruction tuning (Sanh et al., 2022; Wei et al., 2022), in which the finetuning data is accompanied by a prompt containing an instruction. This has been shown to give models the ability to generalize to related tasks and has proven to be effective for MT (Alves et al., 2023, 2024).
@@ 2.2 Robustness to Character Noise
Character perturbations have been shown to have a large negative impact on the performance of MT models (Belinkov and Bisk, 2018). Consequently, a number of techniques have been proposed to mitigate their impact.
@@
A common technique to increase robustness is to train MT models on examples with added source errors. Given that high-quality corpora containing authentic errors are rare, the added noise is generally synthetic (Karpukhin et al., 2019), although it can be tuned to resemble natural errors (Martucci et al., 2021; Vaibhav et al., 2019). Whether training on synthetic noise is actually helpful for becoming robust to natural errors is an open question, with various works coming to contradictory conclusions (Belinkov and Bisk, 2018; Vaibhav et al., 2019). Another hazard is that training on noise can reduce performance on clean data (Khayrallah and Koehn, 2018), so the quantity of noise needs to be chosen carefully.
@@ 3 Robustness to Synthetic Noise
Model | ende | deen | enfr | fren | enko | koen | enpt | pten |
OPUS | 84.61 | 88.24 | 87.74 | 89.21 | 86.60 | 86.35 | 88.98 | 88.55 |
NLLB | 87.82 | 89.31 | 88.11 | 89.25 | 88.51 | 87.59 | 89.23 | 89.54 |
Tower | 88.21 | 89.62 | 88.79 | 89.63 | 89.33 | 88.39 | 89.83 | 89.98 |
GPT-3.5 | 88.52 | 89.64 | 88.83 | 89.45 | 89.04 | 87.98 | 89.83 | 89.81 |
LP | HF Path | Vocab | Params |
ende | opus-mt-en-de | 58.1k | 74.4M |
deen | opus-mt-de-en | 58.1k | 74.4M |
enfr | opus-mt-tc-big-en-fr | 53.0k | 232.7M |
fren | opus-mt-tc-big-fr-en | 53.0k | 232.7M |
enko | - | 32.0k* | 209.2M |
koen | - | 32.0k* | 209.2M |
enpt | opus-mt-tc-big-en-pt | 54.7k | 234.5M |
pten | - | 60.0k | 75.4M |
In our first experiments, we evaluate how various models perform in the presence of token-level synthetic errors. Although synthetic errors differ from “naturally occurring” noise, they are adjustable and function as a stress test for MT systems.
@@ 3.1 Experiments
We use four types of synthetic perturbations, each of which is a plausible error based on the mechanics of typing. For each noise type, we corrupt 10% to 100% of whitespace-delimited tokens in the FLORES-200 devtest data (NLLB Team et al., 2022). We use the following noise types:
@@
chardupe: duplicate a character.
chardrop: delete a character.
key: replace a character with an adjacent one on a keyboard. We use the QWERTZ layout for German, AZERTY for French, QWERTY for English and Portuguese, and South Korean Dubeolsik for Korean.
For Korean, we used hangul-jamo333https://github.com/jonghwanhyeon/hangul-jamo to decompose hangul characters into jamo, which represent individual keystrokes, before applying perturbations.
@@
We use models that differ in their scope (bi- or multilingual), architecture (encoder-decoder or decoder-only), and size (74M-7B parameters).
@@
Tower: We use the 7 billion parameter version of TowerInstruct444https://huggingface.co/Unbabel/TowerInstruct-7B-v0.1 (Alves et al., 2024), a recently-released instruction-tuned LLM derived from Llama 2 (Touvron et al., 2023). It covers translation between 10 languages.
GPT-3.5:555Specifically, we use gpt-3.5-turbo-1106. the architecture and training data of GPT-3.5 are unknown, making it difficult to draw scientific conclusions from its results. We include it because it has shown success at MT (Hendy et al., 2023) and the related GPT-4 has been shown to be able to correct some character perturbations (Cao et al., 2023).
For NLLB, Tower, and the models with a listed HF Path in Table 2, we use public checkpoints from the Hugging Face transformers library.666https://github.com/huggingface/transformers For the pten OPUS model, we use a script777https://github.com/huggingface/transformers/blob/main/src/transformers/models/marian/convert_marian_tatoeba_to_pytorch.py to convert its Marian checkpoint888https://github.com/Helsinki-NLP/Tatoeba-Challenge/tree/master/models/por-eng to transformers format. We directly use checkpoints from the Tatoeba Challenge (Tiedemann, 2020) and perform inference with Marian (Junczys-Dowmunt et al., 2018) for enko999https://object.pouta.csc.fi/Tatoeba-MT-models/eng-kor/opusTCv20210807-sepvoc_transformer-big_2022-07-28.zip and koen101010https://object.pouta.csc.fi/Tatoeba-MT-models/kor-eng/opusTCv20210807-sepvoc_transformer-big_2022-07-28.zip because the conversion script failed for them.
Inference.
We generate translations using beam search (Reddy et al., 1977) with a beam size of 5 for all models except GPT-3.5. For GPT-3.5, we sample with temperature 0. We use prompts for Tower111111‘‘Translate the following text from [source language] to [target language].nSource:[source text]n[target language]:’’ and GPT-3.5.121212‘‘Translate this sentence from [source language] to [target language].nSource:[source text]nTarget:’’
Evaluation.
Our base metric for scoring the translation performance on a corpus is COMET (Rei et al., 2020).131313Specifically, we use COMET-22 (Rei et al., 2022a). COMET computes a normalized score for a hypothesis , conditioned on the source and a reference . When we compute scores for translations from noisy data, we provide the COMET model the clean source, not the noisy version that was actually used to generate hypotheses. We measure the trajectory of performance as the amount of noise is increased, as depicted in Figure 1. To represent this trajectory as a single number, for each configuration we fit a linear regression to predict how much COMET declines relative to the clean performance141414There is no need to learn an intercept term because the decline is relative to the model’s clean performance. as a function of the proportion of noised tokens. We report the learned slope, which we call COMET-slope. The higher (closer to zero) the COMET-slope is, the more robust the model is. This metric can also be interpreted as the number of COMET points that would be lost if every token were corrupted.
@@
-10.27
@@ Results.
The results in Table 1 show that on clean data, Tower and GPT-3.5 are the two strongest systems, with at least one of them recording the highest COMET score for all eight language pairs. The gap between the strongest system and the much smaller OPUS models is at most COMET points for all language pairs except ende. However, Table 3 shows that the differences become much larger on noisy data. For all language pairs and noise types, OPUS suffers far more from perturbations than any of the other models do. On the other end of the spectrum, GPT-3.5 is almost always more robust than other models, sometimes strikingly so. NLLB and Tower are between these two extremes – they are far more robust than OPUS, but only come close to GPT-3.5 for chardupe noise. For swap and chardrop noise, NLLB is more robust than Tower when translating to English, while the reverse is true when translating from English. This trend is less consistent for chardupe noise. For key noise, NLLB is more robust than Tower for every pair except koen.
@@ 3.2 Analysis
Fluency and adequacy.
All models decline as the noise level is increased, but they do not decline in the same way. In Figure 2 we show how the fluency of the model’s predictions, measured by the perplexity of GPT-2 (Radford et al., 2019), is related to the quality of those predictions in terms of COMET. For fren swaps, it is clear that NLLB and GPT-3.5 continue producing fluent English text even as the noise level increases.
@@
From these experiments, it is tempting to conclude that robustness depends largely on model size (OPUS is times smaller than any other system) or on multilinguality (all systems except OPUS are multilingual). However, Figure 3 tells a different story. We reran swap noise experiments with three extra models: the 600M and 1.3 billion parameter versions of NLLB (the former of which is distilled from the 54B version of NLLB) and the 1.2 billion parameter version of M2M (Fan et al., 2021). Despite NLLB-1.3B and M2M-1.2B being multilingual models of similar sizes, they do not respond the same to noise: NLLB-1.3B follows a similar curve to NLLB-3.3B, while M2M-1.2B suffers nearly as much as OPUS. NLLB-600M is somewhere between these extremes. Further work is needed to determine what factors influence this kind of robustness.
@@
Introducing perturbations affects not only translation quality but also runtime. Perturbations create character sequences that are less similar to the data that tokenizers are trained on, which leads to more pieces being used to encode the sentence. This is true even for chardrop noise, which increases the length of the tokenized sequence even as it shortens the detokenized sequence. In Table 4, we compare tokenizers by their fertility — the average number of subword pieces per whitespace word — on clean and key data. While OPUS tokenizers generally have very low fertility on clean data, it increases more than the other tokenizers, suggesting the tokenizer itself is less robust to character perturbations. It is also notable that Tower and GPT-3.5 have high fertility even on clean Korean text. While this is a symptom of tokenizer unfairness in large models (Petrov et al., 2023), it can also be a sign of tokenizer robustness: the higher the fertility, the closer the model is to byte-level tokenization. This results in noisy token sequences that are much closer to the clean sequences for Tower and GPT-3.5, as can be seen in terms of F1 in Table 5. The same trend does not hold for the other languages.
@@
swap | chardrop | chardupe | key | |

### v2

@@ Abstract
Did Translation Models Get More Robust Without Anyone Even Noticing?
Ben Peters and André F. T. Martins Instituto de Telecomunicações, Lisbon, Portugal Instituto Superior Técnico, Universidade de Lisboa, Lisbon, Portugal ELLIS Unit Lisbon (LUMLIS), Lisbon, Portugal Unbabel, Lisbon, Portugal benzurdopeters@gmail.com, andre.t.martins@tecnico.ulisboa.pt
@@ 1 Introduction
For years, the conventional wisdom has been that neural machine translation (MT) models are highly sensitive to source-side artificial and natural noise at inference time (Belinkov and Bisk, 2018). This insight has motivated many works that seek to make MT models more robust to noise through either specialized training (Ebrahimi et al., 2018; Karpukhin et al., 2019; Park et al., 2020; Vaibhav et al., 2019) or bespoke architectures (Rust et al., 2022; Salesky et al., 2021). However, MT is increasingly being performed in a different paradigm than when these analyses and architectures were created. Previously, models were mostly trained from scratch on task-specific data. Nowadays, strong results often depend on instruction-tuned large language models (LLMs) like TowerLLM (Alves et al., 2024) or opaque proprietary systems like ChatGPT.111https://chat.openai.com/ These huge models may make existing robustness techniques more expensive (due to higher parameter counts) or impossible (specialized architectures cannot be grafted onto an existing pretrained system). So the question is, are these robustness techniques still necessary in the era of LLMs, or have larger models and training sets made today’s models sufficiently robust on their own?
In this work, we investigate these questions through experiments on social media text and synthetically noised corpora. These experiments play complementary roles: social media text contains diverse noise phenomena, but their effect is hard to isolate because the errors are unlabeled. On the other hand, synthetic errors differ from real-world noise, but they are interpretable and controllable, offering a way to measure noise in vitro. By evaluating on a broad spectrum of error types, we can paint a more vivid picture of what kinds of noise, and at what quantities, cause problems for MT systems. We make the following contributions:222Our code is available at https://github.com/utter-project/robust-mt.
We show (§3) that large pretrained models are much more robust to synthetic errors than conventional NMT models (see Figure 1), even when they perform similarly on clean data. This result holds across noise types and language pairs, even though the large models lack architectural features that specifically encourage robustness to character noise.
@@
We show (§5) that finetuning on noisy translation data and source correction pipelines are both effective approaches to mitigate synthetic noise without substantially worsening performance on clean data, allowing conventional NMT models to become more robust than GPT-3.5 to 3 out of 4 synthetic noise types. Combining correction with 7-13B parameter LLM-based translation models yields even higher robustness, allowing these pipelines to surpass GPT-3.5 on all of our synthetic benchmarks, often by a wide margin. Although correction is less effective for social media data on the whole, many individual examples benefit from it, suggesting that identifying these examples is a future direction.
@@ 2.1 Architectures for MT
In recent years, mainstream MT techniques have been based on the transformer (Vaswani et al., 2017), which uses multi-headed self-attention to mix information across time steps. In the original work, transformers used an encoder-decoder paradigm similar to recurrent MT models (Bahdanau et al., 2014). These models pair an encoder over the source with a decoder, an autoregressive language model that predicts target tokens one at a time. These tokens usually come from a subword vocabulary (Kudo, 2018; Sennrich et al., 2016). Initially, transformer MT models were trained from scratch for a single language pair on parallel data from sources such as the OPUS parallel corpus collection (Tiedemann, 2012).
@@
Although single language pair models often perform well, they struggle in the absence of large quantities of data, making it difficult to achieve good results in low resource language settings. This problem can be mitigated through multilingual training with systems like M2M-100 (Fan et al., 2021) and NLLB-200 (NLLB Team et al., 2022). Low resource language pairs often benefit from training data in other languages. One challenge is language imbalance – the subword vocabulary and training procedure need to be designed to allow strong performance across covered language pairs in spite of this imbalance.
@@
In parallel to these MT-centric developments, transformers have increasingly been used in a transfer learning set-up in which a model is pretrained on a generic objective for which massive data is available. The model can then be finetuned on one or more downstream tasks. When the pretraining objective is language modeling (Radford et al., 2018), it is straightforward to use the model for generation tasks, including MT (Hendy et al., 2023). Recently, the paradigm has shifted from traditional finetuning to instruction tuning (Sanh et al., 2022; Wei et al., 2022), in which the finetuning data is accompanied by an instructional prompt. This has been shown to give models the ability to generalize to related tasks and has proven effective for MT (Alves et al., 2023, 2024).
@@ 2.2 Robustness to Character Noise
Character perturbations can have a large negative impact on MT model performance (Belinkov and Bisk, 2018). Consequently, a number of techniques have been proposed to mitigate their impact.
@@
A common technique to increase robustness is to train MT models on examples with added source errors. Given that high-quality corpora containing authentic errors are rare, the added noise is generally synthetic (Karpukhin et al., 2019), although it can be tuned to resemble natural errors (Martucci et al., 2021; Vaibhav et al., 2019). Whether training on synthetic noise is actually helpful for becoming robust to natural errors is an open question, with various works coming to contradictory conclusions (Belinkov and Bisk, 2018; Vaibhav et al., 2019).
@@ 3 Robustness to Synthetic Noise
xxen
Model
deen
fren
koen
pten
avg.
OPUS
88.17
89.19
86.35
88.36
88.02
NLLB
89.28
89.29
87.69
89.72
89.00
TI
89.77
89.76
88.69
90.16
89.60
GPT-3.5
89.64
89.45
87.98
89.81
89.22
enxx
Model
ende
enfr
enko
enpt
avg.
OPUS
84.02
87.63
86.58
88.94
86.79
NLLB
88.07
88.30
88.48
89.58
88.61
TI
88.57
89.16
90.12
90.02
89.47
GPT-3.5
88.52
88.83
89.04
89.83
89.05
In our first experiments, we evaluate how models perform in the presence of token-level synthetic errors. Although these errors differ from “naturally occurring” noise, they are adjustable and function as a stress test for MT systems.
@@ 3.1 Experiments
We use four types of synthetic perturbations, each of which is a plausible error based on the mechanics of typing. For each noise type, we create ten noised versions of the FLORES-200 devtest set (NLLB Team et al., 2022) corresponding to noise levels . Within a version of the corpus, each whitespace-delimited token is perturbed with probability and otherwise not altered. Therefore each token can be perturbed at most once. We use the following noise types:
@@
dupe: duplicate a character.
drop: delete a character.
key: replace a character with an adjacent key. Further details are in Appendix A.
@@
We use models that differ in their scope (bi- or multilingual), architecture (encoder-decoder or decoder-only), and size (74M-13B parameters).
@@
TI: We use the 13 billion parameter version of TowerInstruct-v0.1 (Alves et al., 2024), an instruction-tuned LLM that can translate between 10 languages.
As TI and GPT-3.5 were both trained on closed data, it is possible that they were trained on our test sets. We include them because of the lack of high-quality fully-open LLM-based translation systems.
swap
Model
deen
fren
koen
pten
avg.
OPUS
-65.05
-65.41
-36.18
-63.44
-57.52
NLLB
-18.14
-20.97
-23.79
-20.81
-20.93
TI
-27.61
-27.01
-23.45
-25.54
-25.90
GPT-3.5
-4.36
-5.85
-20.89
-6.78
-9.47
drop
Model
deen
fren
koen
pten
avg.
OPUS
-53.61
-49.92
-29.54
-52.48
-46.39
NLLB
-16.50
-17.27
-21.11
-19.04
-18.48
TI
-19.62
-18.01
-17.33
-17.71
-18.17
GPT-3.5
-6.55
-5.68
-17.81
-7.09
-9.28
dupe
Model
deen
fren
koen
pten
avg.
OPUS
-34.31
-31.83
-6.92
-34.58
-26.91
NLLB
-4.07
-5.31
-4.36
-4.58
-4.58
TI
-3.37
-4.54
-1.82
-3.62
-3.38
GPT-3.5
-1.36
-1.42
-5.64
-1.44
-2.47
key
Model
deen
fren
koen
pten
avg.
OPUS
-63.78
-65.40
-38.48
-65.50
-58.29
NLLB
-20.66
-21.86
-28.01
-23.60
-23.53
TI
-29.15
-32.18
-19.80
-34.15
-28.82
GPT-3.5
-9.17
-8.63
-16.31
-10.27
-11.09
@@
-8.26
Inference.
For GPT-3.5, we sample with temperature 0. For other models, we decode with a beam size of 5. Details are in Appendix C.
Evaluation.
Our base corpus-level translation metric is COMET (Rei et al., 2020).444We use Unbabel/wmt22-comet-da (Rei et al., 2022a). COMET computes a normalized score for a hypothesis , conditioned on the source and a reference . When we compute scores for translations from noisy data, we provide the COMET model the clean source, not the noisy version that was actually used to generate hypotheses. We measure the trajectory of performance as the amount of noise is increased, as depicted in Figure 1. To represent this trajectory as a single number, we fit a linear regression to predict the COMET decline relative to the clean score555There is no need to learn an intercept term because the decline is relative to the model’s clean performance. as a function of the proportion of noised tokens. We report the learned slope, which we call COMET-slope. A higher (closer to zero) COMET-slope indicates a more robust model. This metric can be interpreted as the number of COMET points that would be lost if every token were corrupted.
@@ Results.
Table 1 shows that on clean data, TI records the highest COMET for all eight language pairs. The gap between the strongest system and the much smaller OPUS models is less than COMET for all pairs except ende and enko. However, Tables 2 and 3 show that OPUS suffers more from perturbations than the other models do. On the other end of the spectrum, GPT-3.5 is almost always the most robust system. NLLB and TI are between these two extremes. For swap and drop noise, NLLB is more robust than TI when translating to English, while the reverse is true when translating from English. This trend is less consistent for dupe noise. For key noise, NLLB is more robust than TI for every pair except koen. BLEU and chrF results are in Appendix H.
@@ 3.2 Analysis
| Model | clean | swap | drop | dupe | key |
| OPUS | 88.94 | -72.97 | -69.66 | -57.94 | -75.81 |
| OPUSLLM | 85.78 | -73.05 | -69.68 | -55.63 | -74.48 |
@@
From these experiments, one might conclude that robustness depends largely on model size (OPUS is times smaller than any other system) or multilinguality (all except OPUS are multilingual). However, Figure 2 shows that these are not the only factors. We reran swap experiments with NLLB-600M, NLLB-1.3B, and M2M-1.2B (Fan et al., 2021); despite the similar sizes of NLLB-1.3B and M2M-1.2B, and the fact that they are both massively multilingual, they handle noise differently: NLLB-1.3B is similar to NLLB-3.3B, while M2M-1.2B suffers as much as OPUS.
Impact of architecture.
Having shown that model size does not have a strong effect on robustness (at least not for NLLB), we next investigate the impact of architecture on performance. Is the gap between OPUS and other models primarily due to differences in training data, or is there some aspect of the LLMs’ decoder-only structure that encourages robustness? To investigate, we trained a 1.3B parameter666We would have preferred to train a 13B model (similar to TI), but this was impossible due to resource constraints. decoder-only model on the same Tatoeba Challenge data as was used by the enpt OPUS model. Training details are given in Appendix D. The performance and robustness of this model, which we dub OPUSLLM, are shown in Table 4. Although its performance on clean data lags COMET points behind OPUS, the COMET-slope is similar for all four noise types, suggesting that the robustness of recent models is due to their training data, not their size or architecture.
@@
Introducing perturbations affects not only translation quality but also runtime. Perturbations create character sequences that are less similar to the data that tokenizers are trained on, which leads to more pieces being used to encode the sentence. This is true even for drop noise, which increases the length of the tokenized sequence even as it shortens the detokenized sequence. In Table 5, we compare tokenizers by their fertility — the average number of subword pieces per whitespace word — on clean and key data. While OPUS tokenizers generally have very low fertility on clean data, it increases more than the other tokenizers, suggesting the tokenizer itself is less robust to character perturbations. It is also notable that TI and GPT-3.5 have high fertility even on clean Korean text. While this is a symptom of tokenizer unfairness in large models (Petrov et al., 2023), it can also be a sign of tokenizer robustness: the higher the fertility, the closer the model is to byte-level tokenization. This results in noisy token sequences that are much closer to the clean sequences for TI and GPT-3.5, as can be seen in terms of F1 in Table 6. The same trend does not hold for the other languages.
@@
| swap | drop | dupe | key | |